OpenStudy (anonymous):

3. Angelo took a group of people to the movies and spent $80 on tickets. Adult tickets cost $10 and child tickets cost $6. The equation to represent his cost is 10A + 6C = $80. If there were 5 children in the group, how many adult tickets did he purchase? Show your work and use a complete sentence to explain your solution in the context of the original problem.

OpenStudy (anonymous):

questions :What does $80 represent? What does $10 represent? What does $6 represent? What does A represent? What does C represent? What does 5 represent? Which variable can you replace with a number? Solve the equation. Start by replacing one of the variables with a number: 10A + 6C = $80.

OpenStudy (anonymous):

i have all the answers besides solve the equation

OpenStudy (paki):

just do this sir... fir 5 children... just put c=5 in the equation... 10a+6(5)=80 10a+30=80 10a-30+30=80-30 10a=50 10a/10=50/10 a=5 there are 5 children, 30 must be spent, 50 is left, since each adult costs 10, 5 there are 5 adult tickets

OpenStudy (anonymous):

thanks (:

OpenStudy (paki):

pleasure :-)

OpenStudy (anonymous):

can you help me with one more ?

OpenStudy (paki):

sure, i will try...

OpenStudy (anonymous):

4. A quality control inspector is in charge of inspecting drill bits that are produced on an assembly line. The diameter of a drill bit can be modeled by the equation |x − 0.525| = 0.01. Show your work in solving the eqation to determine the minimum and maximum diameters of the drill bits. Solve for the first absolute value Solve for the second absolute value Min: Max:

OpenStudy (paki):

hmmm... see... x−0.525|=0.01 will turns into (x−0.525)=0.01 and −(x−0.525)=0.01

OpenStudy (paki):

then... Solving for x in each equation above gives you -0.515 and 0.535. Assuming x is the diameter, it can not be negative (obviously). So I'm thinking they need to be within 0.515 and 0.535. Don't forget units if there are any.

OpenStudy (paki):

@sataytre

OpenStudy (anonymous):

im confused..

OpenStudy (paki):

why...

OpenStudy (anonymous):

wait nevermind thank you (:

OpenStudy (paki):

pleasure sir... :-)